# Dynamical systems defined by polynomials with algebraic properties

### v4

@@ 5 Stream zeros and the automorphism group
This implies (5). Moreover, let $(c_{0},c_{1},\cdots,c_{k-1})$ be the first row of $B$, then from the above, we have $c_{0}\theta_{i}^{k}+c_{1}\theta_{i}^{k-1}+\cdots+c_{k-1}\theta_{i}=\lambda_{i}\theta_{i}^{k}$, so that
$\lambda_{i}=c_{0}+c_{1}\theta_{i}^{-1}+\cdots+c_{k-1}\theta_{i}^{-(k-1)}~{}(i=1,\cdots,k).$ (6)
Conversely, if $B$ is a $k\times k$ matrix satisfying (5) and (6) with $\lambda_{i}\in\mathbb{C}~{}(i=1,\cdots,k)$ such that $\lambda_{1}\cdots\lambda_{k}=\pm 1$. Then, we have $\det B=\pm 1$. Moreover, since
$M_{P}=\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{-1}&0&\cdots\\ 0&\ddots&0\\ \cdots&0&\theta_{k}^{-1}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)^{-1},$
we have $BM_{P}=M_{P}B,$ since both terms are equal to
$\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)\left(\begin{array}[]{ccc}\lambda_{1}\theta_{1}^{-1}&0&\cdots\\ 0&\ddots&0\\ \cdots&0&\lambda_{k}\theta_{k}^{-1}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)^{-1}.$
Thus, we have $B\in Saut_{P}$. Note that all entries of $B$ are integers since $B$ is a linear combination of
$M_{P}^{l}=\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{-l}&0&\cdots\\ 0&\ddots&0\\ \cdots&0&\theta_{k}^{-l}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)^{-1},$
with integer coefficients $c_{0},c_{1},\cdots,c_{k-1}$ by (6).
(ii) Assume that all the roots of $P(z)=0$ are real. Then, (5) gives an isomorphism from $Saut_{P}$ into
$\{(\lambda_{1},\cdots,\lambda_{k})\in(\mathbb{R}\setminus\{0\})^{k};~{}\lambda_{1}\cdots\lambda_{k}=\pm 1\}.$
If $B$ is mapped to one of them with $\lambda_{i}\neq\pm 1$ for some $i=1,2,\cdots,k$, then it has an infinite order. Otherwise, it is mapped to $(\lambda_{1},\cdots,\lambda_{k})$ with $\lambda_{i}\in\{\pm 1\}~{}(i=1,\cdots,k)$. Hence, the order is either $1$ or $2$.
The final statement follows since $\lambda_{1}$ in (5) determine $B$ and the set of $\lambda_{1}$ of the form $\lambda_{1}=c_{0}+c_{1}\theta_{1}^{-1}+\cdots+c_{k-1}\theta_{1}^{-(k-1)}\notin\{0,\pm 1\}$ with $c_{i}\in\mathbb{Z}~{}(i=0,1,\cdots,k-1)$ is a countable multiplicative subgroup of $\mathbb{R}_{>0}$. ∎
@@ 6 $Saut_{P}$ in the case $\deg P=2,3$
We now give a complete characterization of the elements of $Saut_{P}$ when the degree of polynomial $P(z)$ is two. More precisely, we have
@@
Assume $(3)$ with the degree $k=2$ and discriminant $D=a_{1}^{2}-4a_{2}$.
(i) Assume $D>0$. Then the following conditions are equivalent:
@@
(i-2) $B$ is a $2\times 2$ integer matrix such that $(5)$ holds for some $\lambda_{i}\in\mathbb{C}~{}(i=1,2)$ such that $\lambda_{1}\lambda_{2}=1$,
(i-3)  $B=\left(\begin{array}[]{cc}c_{0}&c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right)$ for some $c_{0},c_{1}\in\mathbb{Z}$ such that
@@
(i-4) $B$ is of the form $\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)\in GL(2,\mathbb{Z})$ satisfying
$\frac{p\theta+p^{\prime}}{q\theta+q^{\prime}}=\theta$ (8)
for some root $\theta$ of the equation $a_{2}z^{2}+a_{1}z+1=0$.
Moreover, $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group if and only if $D$ is a non-square integer. In addition, if $D=4$, then it is a cyclic group of order $2$. Otherwise, it is the trivial group.
@@
(iii) Assume $D<0$. Then $Saut_{P}$ is the trivial group except for the case $D=-4$, where $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order $2$.
@@
The essential part of (ii) of Theorem 6 is to solve the so-called Pell’s equation, i.e., Diophantine equation of the form
@@
It is well known that the main tool to solve it is the theory of continued fraction (see [1, 3], for example).
@@
(i) Assume $D>0$. Let $\theta_{i}~{}(i=1,2)$ be the two roots of $P(z)=0$. By Theorem 5, $B=\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)\in Saut_{p}$ if and only if $pq^{\prime}-p^{\prime}q=\pm 1$ and that
@@
holds for some $\lambda_{i}~{}(i=1,2)$ such that $\lambda_{i}=c_{0}+c_{1}\theta_{i}^{-1}$ with $c_{0},c_{1}\in\mathbb{Z}$. Combing with $\theta_{1}\theta_{2}=\frac{1}{a_{2}},~{}\theta_{1}+\theta_{2}=-\frac{a_{1}}{a_{2}}$, we thus have
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}c_{0}&c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right).$
Furthermore, $pq^{\prime}-p^{\prime}q=\pm 1$ is equivalent to (7). Thus, (i-1) and (i-3) are equivalent. Here, the additional condition that $\lambda_{i}=c_{0}+c_{1}\theta_{i}^{-1}$ with $c_{0},c_{1}\in\mathbb{Z}$ follows immediately if we assume that $B$ is an integer matrix satisfying (10). Hence, they are also equivalent to (i-2).
Let $\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)\in Saut_{p}$. Then by (10), we have
@@
Conversely, assume (i-4). Without loss of generality, assume $\frac{p\theta_{1}+p^{\prime}}{q\theta_{1}+q^{\prime}}=\theta_{1}$. Then since $\theta_{1}$ and $\theta_{2}$ are algebraically conjugate, we also have $\frac{p\theta_{2}+p^{\prime}}{q\theta_{2}+q^{\prime}}=\theta_{2}$. Therefore, there exists $\lambda_{i}~{}(i=1,2)$ such that (10) holds and (i-4) implies (i-2). Thus, the first half of statement (i) is proved.
@@
is an integer matrix. Hence,
$M_{P}B+M_{P}=\frac{1}{\theta_{1}\theta_{2}(\theta_{1}-\theta_{2})}\left(\begin{array}[]{cc}2\theta_{1}\theta_{2}&-2\theta_{1}\theta_{2}^{2}\\ -2\theta_{2}&-2\theta_{2}^{2}\end{array}\right)$
is an integer matrix. Therefore, $\theta_{1}$ and $\theta_{2}$ are rational since they are ratios of its entries. By $\theta_{1},\theta_{2}=\frac{-a_{1}\pm\sqrt{D}}{2}$, this implies that $D$ is a square integer.
Thus, if $D$ is a nonsquare integer, then $Saut_{P}/\{\pm I_{2}\}$ except for the unit contains only elements of infinite order, and later we prove that it does contain elements of infinite order. Let $\eta$ be the mapping defined on $Saut_{P}$ such that for $B\in Saut_{P}$, $\eta(B)=(\lambda_{1},\lambda_{2})$, where $(\lambda_{1},\lambda_{2})$ satisfies (6) with $k=2$. Then $\eta(B)\in(\mathbb{R}\setminus\{0\})^{2}$ and $Image(\eta)$ is a discrete multiplicative subgroup of ($(\mathbb{R}\setminus\{0\})^{2}$. Let $\tilde{\lambda}$ be the smallest element in $\{\lambda_{1}\in(1,\infty);~{}(\lambda_{1},\lambda_{2})\in Image(\eta)\}$. Since $(1,-1)\notin\eta(B)$, just one of $(\tilde{\lambda},\tilde{\lambda}^{-1})$ or $(\tilde{\lambda},-\tilde{\lambda}^{-1})$ is in $Image(\eta)$. This element generates $\tilde{\lambda}$.
Let $D>0$ be a nonsquare integer. We prove that $Saut_{P}/\{\pm I_{2}\}$ contains nontrivial elements using the continued expansion of $\theta$, which is ultimately periodic. Let it be
@@
be the irreducible fractions with nonnegative denominators, where if $k=0$ and the expansion is purely periodic, then define $\frac{C}{G}=\frac{1}{0},~{}\frac{C^{\prime}}{G^{\prime}}=\frac{0}{1}$, and if $k=1$ ($m=1$), then define $\frac{C^{\prime}}{G^{\prime}}=\frac{1}{0}$ ($\frac{E^{\prime}}{F^{\prime}}=\frac{1}{0}$). Finally, define $2\times 2$-matrix $\left(\begin{array}[]{cc}p_{n}&p^{\prime}_{n}\\ q_{n}&q^{\prime}_{n}\end{array}\right)~{}(n=0,1,2\cdots)$ by
@@
Then, they are elements in $Saut_{P}$ different from each other for $n=0,1,\cdots$ so that $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group.
Let $D=4$. Since $D=a_{1}^{2}-4a_{2}$, $a_{1}$ is even, say $a_{1}=2a,a\in\mathbb{Z}$. Then by (10), $(p-ap^{\prime})^{2}-p^{\prime 2}=\pm 1$. Hence, either $p^{\prime}=\pm 1,~{}p=\pm a$ ($\pm$ correspondingly) or $p^{\prime}=0,~{}p=\pm 1$. In the latter case, we have $B=\pm I_{2}$. In the former case, we have $a_{2}=a^{2}-1$, because $D/4=a^{2}-a_{2}=1$. Then by (9),
$B=\pm\left(\begin{array}[]{cc}a&1\\ -a^{2}+1&-a\end{array}\right).$
Note that $\det B=-1$ and $B^{2}=I_{2}$. Thus, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 2.
Let $D>0$ be a square other than 4. By (11), $B\in Saut_{P}$ if and only if $(2p-a_{1}p^{\prime})^{2}-Dp^{\prime 4}=\pm 4$, which is possible only $p=1,~{}p^{\prime}=0$. Thus, $Saut_{P}/\{\pm I_{2}\}$is trivial.

### v5

@@ 6 Stream zeros and the strong automorphism group
Thus, we have (16). Since $B=(b_{ij})_{i,j=1,\cdots,k}$ is an integer matrix and
$b_{i1}\theta_{1}^{k}+b_{i2}\theta_{1}^{k-1}+\cdots+b_{ik}\theta_{1}=\lambda_{1}\theta^{k-i+1}$
holds for any $i=1,\cdots,k$, we have $\lambda_{1}\in\mathbb{Z}\theta_{1}^{i-1}+\mathbb{Z}\theta_{1}^{i-2}+\cdots+\mathbb{Z}\theta_{1}^{i-k}$, and hence (15). From (16), we also have
$\lambda_{j}=b_{i1}+b_{i2}\theta_{j}^{-1}+\cdots+b_{ik}\theta_{j}^{-(k-1)}~~(j=1,2,\cdots,k).$ (17)
Thus, $\lambda_{2},\cdots,\lambda_{k}$ are the conjugates of $\lambda_{1}$ corresponding to $\theta_{2},\cdots,\theta_{k}$. Since $\det B=\pm 1$ and (15), we have $\lambda_{1}\lambda_{2}\cdots\lambda_{k}=\pm 1$, which implies that the norm of $\lambda_{1}$ is $\pm 1$, and hence, $\lambda_{1}$ is a unit.
Conversely, if $\lambda_{1}$ is a unit satisfying (15). Define $B$ with the cojugates $\lambda_{2},\cdots,\lambda_{k}$ of $\lambda_{1}$ by (16). Then, $B$ is a integer matrix since (15), (17) and the representations as (17) are unique. Moreover it commutes with $C_{P}$ and $\det B=\pm 1$, we have $B\in Saut_{P}$. Thus, we have a bijection bewtween $Saut_{P}$ and the set of units in $\mathbb{Q}(\theta_{1})$ satisfying (15). It is clear that this is an isomorphism with respect to the multiplications.
(ii) By the unit theorem of Dirichlet (Boreivich and Shafarevich [1]), the set of units in $\mathbb{Q}(\theta_{1})$ is a direct product of $c_{1}+c_{2}$ number of cyclic groups of infinite orders with a finite group, where $k_{1},2k_{2}$ are the numbers of real roots and complex roots of $P(z)=0$, respectively. Since the set of units satisfying (15) is a subpgroup of it which is isomorphic to $Saut_{P}$, $Saut_{P}$ is a direct product of $l$ number of cyclic groups of infinite orders with a finite group, where $l\leq k_{1}+k_{2}-1$. ∎
Example 3.
Let $P(z)=2z^{2}-2z-3$. It has the roots $(1\pm\sqrt{7})/2$. Let $\theta_{1}=(1+\sqrt{7})/2$. The fundamental units in $\mathbb{Q}(\sqrt{7})$ are $8\pm 3\sqrt{7}$. Let $\alpha=8+3\sqrt{7}$. Then, it satisfies (15) since
$\alpha=6\theta_{1}+5=9\theta_{1}^{-1}+11.$
Let $B\in Saut_{P}$ be the matrix corresponding to $\lambda_{1}=\alpha$ in (15). Then we have $B=\begin{pmatrix}11&9\\ 6&5\end{pmatrix}$. Hence, $Saut_{P}/\pm I_{2}=\left\{\begin{pmatrix}11&9\\ 6&5\end{pmatrix}^{n};~n\in\mathbb{Z}\right\}$ follows.
Example 4.
Let $P(z)=z^{2}-125$. It has the roots $\pm 5\sqrt{5}$ and the fundamental units in $\mathbb{Q}(\sqrt{5})$ are $\frac{1\pm\sqrt{5}}{2}$. Let $\alpha=\frac{1+\sqrt{5}}{2}$. The smallest power $k$ such that $\alpha^{k}$ satisfies (15) is 8. Let $\lambda_{1}:=\alpha^{8}=682+305\sqrt{5}$. Let $B\in Saut_{P}$ be the matrix corresponding to $\lambda_{1}$ in (15). Then we have $B=\begin{pmatrix}682&7625\\ 61&682\end{pmatrix}$. Hence,
$Saut_{P}/\pm I_{2}=\left\{\begin{pmatrix}682&7625\\ 61&682\end{pmatrix}^{n};n\in\mathbb{Z}\right\}.$
Example 5.
Let $\theta_{1},\theta_{2},\theta_{3}$ be the roots of $P(z)=2z^{3}-3z^{2}-3z+3$ which are $3$ distinct real numbers. Then, we have
$C_{P}=\begin{pmatrix}0&1&0\\ 0&0&1\\ -2/3&1&1\end{pmatrix}$.
We have $\mathbb{Z}[3\alpha]\subsetneq\mathbb{O}_{K}=\mathbb{Z}+3\alpha\mathbb{Z}+3\alpha^{2}\mathbb{Z}$ and $\mathbb{O}_{K}^{*}=\{\pm\epsilon_{1}^{n}\epsilon_{2}^{m}\ |\ m,n\in\mathbb{Z}\}$ with
$\epsilon_{1}=1-3\alpha^{2},\ \epsilon_{2}=5-6\alpha^{2}$
where $\epsilon_{1},\epsilon_{2}\in\mathbb{O}_{K}^{*}$ and $\epsilon_{1},\epsilon_{2}\not\in\mathbb{Z}[3\alpha]$. In this case, we have
$B_{1}=\iota(\epsilon_{1})=E_{3}-3C_{P}^{2}=\begin{pmatrix}1&0&-3\\ 2&-2&-3\\ 2&-1&-5\\ \end{pmatrix}$,
$B_{2}=\iota(\epsilon_{2})=5E_{3}-6C_{P}^{2}=\begin{pmatrix}5&0&-6\\ 4&-1&-6\\ 4&-2&-7\\ \end{pmatrix}$
and both of them are in $M_{3\times 3}(\mathbb{Z})$. Therefore
$Saut=\{\pm B_{1}^{n}B_{2}^{m}\ |\ m,n\in\mathbb{Z}\}.$
Considering the inverse $1/\alpha$, from regular representation by $\{1,1/\alpha,1/\alpha^{2}\}$, we have
$\tilde{C}_{P}=\begin{pmatrix}0&1&0\\ 0&0&1\\ -3/2&3/2&3/2\\ \end{pmatrix}$.
We have $\mathbb{Z}[2/\alpha]\subsetneq\mathbb{O}_{K}=\mathbb{Z}+2/\alpha\mathbb{Z}+(1/\alpha+2/\alpha^{2})\mathbb{Z}$ and
$\epsilon_{1}=-5-1/\alpha+2/\alpha^{2},\ \epsilon_{2}=-7-2/\alpha+4/\alpha^{2}$
and notice that $\epsilon_{2}\in\mathbb{Z}[2/\alpha]$. Thus
$\tilde{B}_{1}=\iota(\epsilon_{1})=-5E_{3}-\tilde{C}_{P}+2\tilde{C}_{P}^{2}=\begin{pmatrix}-5&-1&2\\ -3&-2&2\\ -3&0&1\\ \end{pmatrix}$,
$\tilde{B}_{2}=\iota(\epsilon_{2})=-7E_{3}-2\tilde{C}_{P}+4\tilde{C}_{P}^{2}=\begin{pmatrix}-7&-2&4\\ -6&-1&4\\ -6&0&5\\ \end{pmatrix}$
and both of them are in $M_{3\times 3}(\mathbb{Z})$. Therefore
$\widetilde{Saut}=\{\pm\tilde{B}_{1}^{n}\tilde{B}_{2}^{m}\ |\ m,n\in\mathbb{Z}\}.$
Example 6.
Let $\alpha$ by a root of $2x^{3}-3x^{2}+2x+3$ with $r_{1}=1$ and $r_{2}=1$. Similarly we have
$C_{P}=\begin{pmatrix}0&1&0\\ 0&0&1\\ -3/2&-1&3/2\\ \end{pmatrix}$.
We have $\mathbb{Z}[2\alpha]\subsetneq\mathbb{O}_{K}=\mathbb{Z}+2\alpha\mathbb{Z}+(1/2+3\alpha/2+\alpha^{2})\mathbb{Z}$ and $\mathbb{O}_{K}^{*}=\{\pm\epsilon^{n}\ |\ n\in\mathbb{Z}\}$ with
$\epsilon=\frac{1}{2}+\frac{1}{2}\alpha-\alpha^{2}$ (18)
and $\epsilon\not\in\mathbb{Z}[2\alpha]$. Let
$V=\iota(\epsilon)=\frac{1}{2}E_{3}+\frac{1}{2}C_{P}-C_{P}^{2}=\begin{pmatrix}1/2&1/2&-1\\ 3/2&3/2&-1\\ 3/2&5/2&0\\ \end{pmatrix}$.
Then we see that
$B=V^{3}=\begin{pmatrix}-4&-5&2\\ -3&-6&-2\\ 3&-1&-9\\ \end{pmatrix}$
belongs to $M_{3\times 3}(\mathbb{Z})$ and we have
$Saut=\{\pm B^{n}\ |\ n\in\mathbb{Z}\}.$
We see $\epsilon^{3}=-4-5\alpha+2\alpha^{2}\in\mathbb{Z}[\alpha]\setminus\mathbb{Z}[2\alpha]$. We claim that $\epsilon$ and $\epsilon^{2}$ do not belongs to $\mathbb{Z}[\alpha]$. Assume that $\epsilon\in\mathbb{Z}[\alpha]$ and take an expression $\epsilon=\sum_{i=0}^{\ell}x_{i}\alpha^{i}$ with $x_{i}\in\mathbb{Z}$ with the minimum $\ell$. From (18), we know that $\ell\geq 3$. Then $\rm{Tr}(C_{P}^{n})\ (n=0,1,\dots)$ form a sequence
[…]
which satisfies the linear recurrence
$a_{n+3}=\frac{3}{2}a_{n+2}-a_{n+1}-\frac{3}{2}a_{n}$
and we can show that its denominator is $2^{n}\ (n=0,1,\dots)$. Since
$2=\rm{Tr}V=\sum_{i=0}^{\ell}x_{i}\rm{Tr}(C_{P}^{i})$
we know that $x_{\ell}$ is even. However then using
$\alpha^{\ell}=\frac{3}{2}\alpha^{\ell-1}-\alpha^{\ell-2}-\frac{3}{2}\alpha^{\ell-3}$
we have an expression of $\epsilon$ in $\mathbb{Z}[\alpha]$ of length $\ell-1$, which is a contradiction. Similarly we see $\epsilon^{2}\not\in\mathbb{Z}[\alpha]$. The claim is shown.
Example 7.
Let $\alpha$ be a root of $3x^{4}-2x^{3}+x^{2}+x-2$ with $r_{1}=2$ and $r_{2}=1$. Then
$C_{P}=\begin{pmatrix}0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ 2/3&-1/3&-1/3&2/3\\ \end{pmatrix}$.
We have $\mathbb{O}_{K}=\mathbb{Z}+3\alpha\mathbb{Z}+(\alpha+3\alpha^{2})\mathbb{Z}+(2\alpha+\alpha^{2}+3\alpha^{3})\mathbb{Z}$ and
$\mathbb{O}_{K}^{*}=\{\pm\epsilon_{1}^{m}\epsilon_{2}^{n}\ |\ m,n\in\mathbb{Z}\}$
with
$\epsilon_{1}=1-2\alpha-2\alpha^{2}+3\alpha^{3},\ \epsilon_{2}=361+237\alpha+93\alpha^{2}+468\alpha^{3}.$
In this case, we have
$B_{1}=\iota(\epsilon_{1})=\begin{pmatrix}1&-2&-2&3\\ 2&0&-3&0\\ 0&2&0&-3\\ -2&1&3&-2\\ \end{pmatrix}$,
$B_{2}=\iota(\epsilon_{2})=\begin{pmatrix}361&237&93&468\\ 312&205&81&405\\ 270&177&70&351\\ 234&153&60&304\\ \end{pmatrix}$
and
$Saut=\{\pm B_{1}^{n}B_{2}^{m}\ |\ m,n\in\mathbb{Z}\}.$
@@ 7 $Saut_{P}$ in the case $\deg P=2,3$
[…]
(i) Assume $D>0$ and let the roots of $P(z)=0$ be $\theta_{1},\theta_{2}$. Then the following conditions for an integer matrix $B=\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)$ are equivalent:
@@
[…]
(i-3)  $B=\left(\begin{array}[]{cc}c_{0}&a_{0}c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right)$ for some $c_{0},c_{1}\in\mathbb{Z}$ such that
@@
(i-4) $B$ satisfies that $pq^{\prime}-p^{\prime}q=\pm 1$ and
$\frac{p\theta_{i}+p^{\prime}}{q\theta_{i}+q^{\prime}}=\theta_{i}~~(i=1,2).$ (20)
Moreover, $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group if and only if $D$ is a non-square integer. In addition, if $D\in\{1,4\}$, then it is a cyclic group of order $2$. Otherwise, it is the trivial group.
@@
(iii) Assume $D<0$. Then $Saut_{P}/\{\pm I_{2}\}$ is the trivial group except for the case $D\in\{-3,-4\}$. If D=3D=-3, then $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order $3$. If D=4D=-4, then $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order $2$.
@@
The equation (19) is known as the Pell’s equation, i.e., Diophantine equation of the form
@@
The main tool to solve it is the theory of continued fraction (see [2, 7], for example).
@@
(i) Assume $D>0$.
By Theorem 7, we already know that (i-1) and (i-2) are equivalent.
Assume (i-2). Then we have
@@
with some $\lambda_{i}~(i=1,2)$ such that $\lambda_{1}\lambda_{2}=\pm 1$. It follows from (22) that $\lambda_{i}=c_{0}+c^{\prime}_{1}\theta_{i}^{-1}$ with $c_{0}=p,~c^{\prime}_{1}=p^{\prime}$. Combining with $\theta_{1}\theta_{2}=\frac{a_{0}}{a_{2}},~\theta_{1}+\theta_{2}=-\frac{a_{1}}{a_{2}}$, we thus have
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}c_{0}&c^{\prime}_{1}\\ -\frac{a_{2}}{a_{0}}c^{\prime}_{1}&c_{0}-\frac{a_{1}}{a_{0}}c^{\prime}_{1}\end{array}\right).$
As $B$ is a integer matrix, $\frac{a_{2}}{a_{0}}c^{\prime}_{1}$ and $\frac{a_{1}}{a_{0}}c^{\prime}_{1}$ are integers. Since $a_{2}z^{2}+a_{1}z+a_{0}$ is a primitive polynomial, this is possible only if $c^{\prime}_{1}$ is a multiple of $a_{0}$. Putting $c_{1}=\frac{c^{\prime}_{1}}{a_{0}}\in\mathbb{Z}$, we have $c_{1}^{\prime}=a_{0}c_{1}$ and
$B=\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}c_{0}&a_{0}c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right).$
Clearly, $pq^{\prime}-p^{\prime}q=\pm 1$ is equivalent to (19). Thus, (i-2) implies (i-3).
Conversely, if $B=\left(\begin{array}[]{cc}c_{0}&a_{0}c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right)$, then $BC_{P}=C_{P}B$ follows. Thus, (i-3) implies (i-2).
Let $B\in Saut_{p}$. Then by (16), we have
@@
Assume (i-4). Then, therer exists $\lambda_{1},\lambda_{2}$ such that
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)\left(\begin{array}[]{c}\theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\lambda_{i}\left(\begin{array}[]{c}\theta_{i}^{2}\\ \theta_{i}\end{array}\right)~~(i=1,2).$
Hence, we have (16) and $\lambda_{1}\lambda_{2}=pq^{\prime}-p^{\prime}q=\pm 1$. Thus, (i-4) implies (i-2) and the first half of statement (i) is proved.
@@
$=\frac{1}{\theta_{1}-\theta_{2}}\left(\begin{array}[]{cc}\theta_{1}+\theta_{2}&-2\theta_{1}\theta_{2}\\ 2&-(\theta_{1}+\theta_{2})\end{array}\right)=\frac{1}{\theta_{1}-\theta_{2}}\left(\begin{array}[]{cc}-\frac{a_{1}}{a_{2}}&-\frac{2a_{0}}{a_{2}}\\ 2&\frac{a_{1}}{a_{2}}\end{array}\right)$
is an integer matrix. Therefore $\theta_{1}-\theta_{2}$ is rational, and hence, $D$ is a square integer.
Thus, if $D$ is a nonsquare integer, then $Saut_{P}/\{\pm I_{2}\}$ except for the unit contains only elements of infinite order, and later we prove that it does contain elements of infinite order. Let $\eta$ be the mapping defined on $Saut_{P}$ such that for $B\in Saut_{P}$, $\eta(B)=(\lambda_{1},\lambda_{2})$, where $(\lambda_{1},\lambda_{2})$ satisfies (22). Then $\eta(B)\in(\mathbb{R}\setminus\{0\})^{2}$ and $Image(\eta)$ is a discrete multiplicative subgroup of $(\mathbb{R}\setminus\{0\})^{2}$. Let $\tilde{\lambda}$ be the smallest element in $\{\lambda_{1}\in(1,\infty);~(\lambda_{1},\lambda_{2})\in Image(\eta)\}$. Since $(1,-1)\notin\eta(B)$, just one of $(\tilde{\lambda},\tilde{\lambda}^{-1})$ or $(\tilde{\lambda},-\tilde{\lambda}^{-1})$ is in $Image(\eta)$. This element generates $\tilde{\lambda}$.
Let $D>0$ be a nonsquare integer. We prove that $Saut_{P}/\{\pm I_{2}\}$ contains nontrivial elements using the continued fraction expansion of $\theta_{1}$, which is eventually periodic. Let it be
@@
are irreducible fractions with nonnegative denominators, where if $k=0$, that is, the continued fraction expansion is purely periodic, then we put
$\left(\begin{array}[]{cc}C&C^{\prime}\\ G&G^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$
and if $k=1$ or $m=1$, we put
$\left(\begin{array}[]{cc}C&C^{\prime}\\ G&G^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}c_{0}&1\\ 1&0\end{array}\right)~\mbox{and}~\left(\begin{array}[]{cc}E&E^{\prime}\\ F&F^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}e_{0}&1\\ 1&0\end{array}\right).$
Finally, define $2\times 2$-matrix $\left(\begin{array}[]{cc}p_{n}&p^{\prime}_{n}\\ q_{n}&q^{\prime}_{n}\end{array}\right)~(n\in\mathbb{Z})$ by
@@
Then, they are different each other and the totality of elements in $Saut_{P}/\{\pm I_{2}\}$. Hence, $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group.
Let $D>0$ be a square. By (i-3), for an integer matrix $B=\left(\begin{array}[]{cc}c_{0}&a_{0}c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right)$, $B\in Saut_{P}$ if and only if for some $c_{0},c_{1}\in\mathbb{Z}$ such that (19) holds, That is,
$(2c_{0}-a_{1}c_{1})^{2}-Dc_{1}^{2}=\pm 4.$
It always has solution $c_{0}=\pm 1,~c_{1}=0$ corresponding to $B=\pm I$. The other solutions come from the case $c_{1}\neq 0$. Since the difference between 2 square numbers becomes 4 only if one is 0 and the other is 4, the solution of (19) with $c_{1}\neq 0$ is limited to the following 2 cases:
$\left\{\begin{array}[]{c}D=1\\ c_{1}=\pm 2\\ a_{1}=\pm c_{0}\end{array}\right.(\pm~\mbox{correspondinfly}),~\mbox{and}~\left\{\begin{array}[]{c}D=4\\ c_{1}=\pm 1\\ a_{1}=\pm 2c_{0}\end{array}\right.(\pm~\mbox{correspondinfly})$
In the first case, we have $B=\pm\left(\begin{array}[]{cc}a_{1}&2a_{0}\\ -2a_{2}&-a_{1}\end{array}\right)$ which satisfies that $B^{2}=I_{2}$ in vertue of $a_{1}^{2}-4a_{0}a_{2}=1$. Thus, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 2.
In the second case, note that $a_{1}$ is even and we have $B=\pm\left(\begin{array}[]{cc}\frac{a_{1}}{2}&a_{0}\\ -a_{2}&-\frac{a_{1}}{2}\end{array}\right)$ which satisfies that $B^{2}=I_{2}$ in vertue of $a_{1}^{2}-4a_{0}a_{2}=4$. Thus, Thus, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 2.